### v1

@@ 4.1 Data Augmentation
Swap: As disorder of words between adjacent words dose not alter its overall meaning significantly, the switch of words can be regarded as adding noise to the context. In our approach, we randomly select a span from the context and a position within span, then interchange adjacent word sets.
@@ 5.4 Results
From the perspective of how different model handles the situation of noise, GPT3.5-Turbo outperforms others in every scenarios except when confronted with conflicting contexts (MSConf). Despite ChatGLM’s general inferiority to other models, its capability of managing conflict is surprisingly noteworthy.
@@
MsCons Text:…Mitchell Tower, for example, is modeled after Oxford’s Magdalen Tower… Triples: …, Mitchell Tower ||| modeled after ||| Oxford’s Magdalen Tower, university Commons, … GPT3.5-Turbo: magdalen tower
@@
MsConf Text:…Mitchell Tower, for example, is modeled after Oxford’s Magdalen Tower… Triples: … Mitchell Tower ||| modeled after ||| Oxford’s Radcliffe Camera, university Commons, … GPT3.5-Turbo: oxford’s Radcliffe Camera
@@
Case study To give a more intuitive visualization, some qualitative comparison results are shown in Table 2. It is observed that for SS, MsCons and MsIncons samples, the listed models provide correct answers consistently. However, for SSIncomp sample, while GPT3.5-Turbo is capable of generating a correct response based on its internal knowledge, Baichuan2-13B-Chat fails to produce a specific answer. Although fine-tuning reinforces model’s confidence on the incorrect answers, BC2-13B-SFT-Contrast, guided by further contrastive learning, appropriately declines to answer when lacking internal knowledge. As for MsConf sample, both Baichuan2-13B-Chat and GPT3.5-Turbo are misled by incorrect triples, whereas fine-tuned models exhibits robustness against such interference.

### v3

@@ 4.1 Data Augmentation
Swap: As disorder of words between adjacent words does not alter its overall meaning significantly, the switch of words can be regarded as adding noise to the context. In our approach, we randomly select a span from the context and a position within span, then interchange adjacent word sets.
@@ 5.4 Results
From the perspective of how different model handles the situation of noise, GPT3.5-Turbo outperforms others in every scenario except when confronted with conflicting contexts (MSConf). Despite ChatGLM’s general inferiority to other models, its capability of managing conflict is surprisingly noteworthy.
@@
MS-Cons Text:…Mitchell Tower, for example, is modeled after Oxford’s Magdalen Tower… Triples: …, Mitchell Tower ||| modeled after ||| Oxford’s Magdalen Tower, university Commons, … GPT3.5-Turbo: magdalen tower
@@
MS-Conf Text:…Mitchell Tower, for example, is modeled after Oxford’s Magdalen Tower… Triples: … Mitchell Tower ||| modeled after ||| Oxford’s Radcliffe Camera, university Commons, … GPT3.5-Turbo: oxford’s Radcliffe Camera
@@
Case study To give a more intuitive visualization, some qualitative comparison results are shown in Table 2. It is observed that for SS, MSCons and MSIncons samples, the listed models provide correct answers consistently. However, for SSIncomp sample, while GPT3.5-Turbo is capable of generating a correct response based on its internal knowledge, Baichuan2-13B-Chat fails to produce a specific answer. Though fine-tuning reinforces model’s confidence on the incorrect answers, BC2-13B-SFT-Contrast, guided by further contrastive learning, appropriately declines to answer when lacking internal knowledge. As for MSConf sample, both Baichuan2-13B-Chat and GPT3.5-Turbo are misled by incorrect triples, whereas fine-tuned models exhibits robustness against such interference.